OpenStudy (anonymous):

y^-1+x^-1/(xy)^-1. Explain steps.

myininaya (myininaya):

\[\frac{y^{-1}+x^{-1}}{(xy)^{-1}}=\frac{\frac{1}{y}+\frac{1}{x}}{\frac{1}{xy}}=\frac{\frac{x+y}{xy}}{\frac{1}{xy}}=x+y\]